# Diffusion Models, Image Super-Resolution And Everything: A Survey

### 01 Jan 2024 v1

@@ Abstract
Diffusion Models (DMs) represent a significant advancement in image Super-Resolution (SR), aligning technical image quality more closely with human preferences and expanding SR applications. DMs address critical limitations of previous methods, enhancing overall realism and details in SR images. However, DMs suffer from color-shifting issues, and their high computational costs call for efficient sampling alternatives, underscoring the challenge of balancing computational efficiency and image quality. This survey gives an overview of DMs applied to image SR and offers a detailed analysis that underscores the unique characteristics and methodologies within this domain, distinct from broader existing reviews in the field. It presents a unified view of DM fundamentals and explores research directions, including alternative input domains, conditioning strategies, guidance, corruption spaces, and zero-shot methods. This survey provides insights into the evolution of image SR with DMs, addressing current trends, challenges, and future directions in this rapidly evolving field.
@@
Diffusion Models, Image Super-Resolution, Artificial Intelligence, Survey.
1 Introduction
In the ever-evolving field of computer vision, Super-Resolution (SR) – the enhancement of Low-Resolution (LR) images into High-Resolution (HR) ones – stands as a longstanding yet still perplexing challenge, owing to its inherently ill-posed nature: Several HR images can be valid for any given LR image due to many aspects like brightness and coloring [1]. Spanning applications from natural [2, 3] to sophisticated satellite [4] and medical imaging [5], SR’s progress has been fueled by rapid strides in deep learning.
Diffusion Models (DMs) have recently risen to prominence as a leading family of generative models, challenging the longstanding dominance of Generative Adversarial Networks (GANs) [6, 7, 8, 9]. While earlier generative models showcased impressive image-generation capabilities, they carried inherent limitations. For instance, Autoregressive models can be prohibitively costly for generating HR images [10, 11, 12]. On the other hand, NFs and VAEs often produce samples of sub-optimal quality. Furthermore, GANs demand carefully designed regularization and optimization strategies to manage optimization instability and prevent mode collapse [13]. Nevertheless, these models collectively contributed significantly to the field of SR.
The advent of DMs heralded a new era of generative models and profoundly impacted the realm of generative AI, reinvigorating the field of image SR once again. However, as the volume of research on DMs continues to expand, it is becoming more challenging, particularly for those new to the field, to stay updated on the latest developments. The sheer breadth of this research area can obscure overarching trends and potentially hinder the progression of further research. We tackle these issues by delivering a comprehensive overview of the current research on DMs in image SR.
This survey builds upon our previous work Hitchhiker’s Guide to Super-Resolution [14], which gives a broad overview of the field of image SR in general. We aim to provide a valuable entry point for those new to the domain of DMs in image SR while offering a broader perspective to those already experienced.
@@ 1 Introduction
Section 2 - Super-Resolution Basics: This section provides fundamental definitions and introduces standard datasets, methods, and metrics for assessing image quality commonly utilized in image SR publications.
Section 3 - Diffusion Models Basics: Introduces the principles and various formulations of DMs, including Denoising Diffusion Probabilistic Models (DDPMs), Score-based Generative Models (SGMs), and Stochastic Differential Equations (SDEs). This section also explores how DMs relate to other generative models.
Section 4 - Improvements for Diffusion Models: Common practices for enhancing DMs, focusing on efficient sampling techniques and improved likelihood estimation.
Section 5 - Diffusion Models for Image SR: Presents concrete realizations of DMs in SR, explores alternative domains (latent space and wavelet domain), discusses architectural designs and multiple tasks with Null-Space Models, and examines alternative corruption spaces.
Section 6 - Domain-Specific Applications: DM-based SR applications, namely medical imaging, blind face restoration, atmospheric turbulence in face SR, and remote sensing.
Section 7 - Discussion and Future Work: Common problems of DMs for image SR and noteworthy research avenues for DMs specific to image SR.
Section 8 - Conclusion: Summarizes the survey.
2 Super-Resolution Basics
Image Super-Resolution (SR) represents a collection of methodologies designed to construct High-Resolution (HR) images from one or more Low-Resolution (LR) images. The domain of SR can be broadly categorized into two areas [14]: Single Image Super-Resolution (SISR) and Multi-Image Super-Resolution (MISR). In the SISR paradigm, a single LR image is processed to yield a corresponding HR image. In contrast, MISR methods utilize multiple LR images to synthesize one or many HR outputs. This section primarily focuses on the SISR facet of SR for the context of DMs, exploring relevant datasets, established SR models, and techniques employed to assess image quality.
2.1 Single Image Super-resolution (SISR)
Single Image Super-Resolution (SISR) defines the recovery of a High-Resolution (HR) image from a Low-Resolution (LR) image. Given a LR image $\mathbf{x}\in\mathbb{R}^{\bar{w}\times\bar{h}\times c}$, the objective is to produce a corresponding HR image $\mathbf{y}\in\mathbb{R}^{w\times h\times c}$, where $\bar{w}<w$ and $\bar{h}<h$. This process is often represented through a degradation mapping, defined as
@@ 2.1 Single Image Super-resolution (SISR)
where $\mathcal{D}$ is the degradation map $\mathcal{D}:\mathbb{R}^{w\times h\times c}\rightarrow\mathbb{R}^{\bar{w}\times\bar{h}\times c}$ and $\Theta$ signifies the degradation parameters, encompassing aspects such as blur type, noise, scaling factor, among others. The degradation process is typically uncertain, posing the challenge for SISR to discern the inverse mapping of $\mathcal{D}$. The inherent complexity of this task arises from the ill-posed nature of predicting $\Theta$, leading to an optimization task aimed at minimizing the discrepancy between the predicted SR image $\mathbf{\hat{y}}$ and the original HR image $\mathbf{y}$:
@@
2.2 Datasets
Several datasets, offering a variety of images, resolutions, and content types, are available for SISR and instrumental in effectively training and evaluating SR models. Typically, these datasets consist of LR and HR image pairs. However, certain datasets contain only HR images, with the pairs created by downsampling them using techniques such as bicubic downsampling with anti-aliasing - a default setting for imresize in MATLAB [15].
For natural images, commonly employed datasets include the Berkeley Segmentation Dataset (BSDS300) [3], Flickr2K [16], and the Diverse 2K resolution (DIV2K) dataset [17]. Additionally, datasets such as ImageNet [18] and the Visual Object Classes (VOC2012) [19], which are favored in the Computer Vision (CV) community, are also utilized for SISR tasks. In face SR, datasets featuring high-quality, diverse images of human faces are required. To meet this demand, datasets like CelebA-HQ [20] and Flickr-Faces-HQ (FFHQ) [21] are commonly employed.
2.3 SR Models
@@
where $\mathbf{\hat{y}}$ is the predicted HR approximation of the LR image $\mathbf{x}$ and $\theta$ the parameters of $\mathcal{M}$. Consequently, this translates into an optimization objective that minimizes Equation 2, the difference between the estimation $\mathbf{\hat{y}}$ and the ground-truth HR image $\mathbf{y}$ under a given loss function $\mathcal{L}$. The following sections will focus on standard methods to develop such a SR model, especially conventional deep learning methods, before we examine in detail how diffusion models fulfill the role of a SR model. This includes a brief overview of conventional methods and a short examination of state-of-the-art deep learning methods.
2.3.1 Traditional Methods
Traditional methods for image SR define a range of methodologies, such as statistical [22], edge-based [23], [24] patch-based [25], [26], prediction-based [27], [28] and sparse representation techniques [29]. These methodologies fundamentally rely on image statistics and the information inherent in existing pixels to generate HR images. Despite their utility, a noteworthy drawback of these traditional methods is the potential introduction of noise, blur, and visual artifacts.
2.3.2 Deep Learning Methods
With advancements in Deep Learning (DL) and computational power, image SR has significantly evolved, with DL-based approaches providing notable improvements over traditional methods. DL-based SR methods mainly employ Convolutional Neural Networks (CNNs) for end-to-end mapping from LR to HR images. Initial models, such as SRCNN [30], FSRCNN [31], and ESPCNN [32], utilized simple CNNs of diverse depth and feature maps sizes.
Subsequently, concepts from the broader CV domain were incorporated into SR models, e.g., ResNet led to SRResNet [33], where residual information was propagated to successive network layers. Two approaches were introduced to learn residuals - local and global, each differing in their objectives and internal network connections. DenseNet [34] was also adapted to this task through the use of dense blocks where each layer receives the feature maps generated in all the preceding layers in addition to the inputs like in SRDenseNet [35] and ESRGAN [36]. Recursive CNNs that recursively use the same module to learn representations were also inspired by other CV methods for DL-based SR methods in DRCN [37], DRRN [38] and CARN [39]. More recently, attention mechanisms have been incorporated to enhance feature extraction by focusing on regions of interest in images, predominantly via the channel and spatial attention mechanisms [14].
@@ 2.3.2 Deep Learning Methods
2.3.3 Pixel-based Methods
Pixel-based techniques utilize loss functions designed to minimize the difference directly in pixel space. Commonly used loss functions are the L1 and L2 losses. Models trained with pixel-based loss functions tend to generate images with a tendency toward blurriness and excessive smoothing, which can result in less high-frequency details and even lead to unrealistic visual appearances [14]. One way to overcome this limitation is to use generative models, such as GANs, flow-based methods, or diffusion models.
2.3.4 Generative Adversarial Networks (GANs)
@@
2.3.5 Flow-based Methods
@@
2.4 Image Quality Assessment (IQA)
@@
2.4.1 Peak Signal-to-Noise Ratio (PSNR)
@@
2.4.2 Structural Similarity Index (SSIM)
@@
2.4.3 Mean Opinion Score (MOS)
The MOS is a subjective measure that leverages human perceptual quality for the evaluation of the generated SR images. Human viewers are shown SR images and asked to rate them with quality scores that are then mapped to numerical values and later averaged. Typically, these range from 1 (bad) to 5 (good) but may vary. While this method is a direct evaluation of human perception, it is more time-consuming and cumbersome to conduct compared to objective metrics. Moreover, due to the highly subjective nature of this metric, it is susceptible to bias.
2.4.4 Consistency
@@
2.4.5 Learned Perceptual Image Patch Similarity (LPIPS)
@@
2.4.6 Fool Rate
The Fool Rate is a 2-Alternative Forced-Choice (2AFC) paradigm that measures how humans discriminate true images from model-generated ones using two tasks [42]. In the first task, the subjects are asked to pick the HR image they believe is of higher quality. They are shown an LR input between two HR images of the ground truth and the generated SR image. In task 2, they are asked to guess which image they believe to be from a camera between the HR ground truth and generated SR image. The fool rate is the fraction of trials in which the subjects chose the generated SR over the ground truth.
Figure 1: Principle of Diffusion Models. The forward diffusion adds noise iteratively (red). The backward diffusion, the iterative refinement process, reverts the process (blue).
3 Diffusion Models Basics
Diffusion Models (DMs) have profoundly impacted the realm of generative AI, and many approaches that fall under the umbrella term DM have emerged. At their core, these models share a fundamental principle called diffusion: generate data through the gradual introduction of noise into an initial, often simplistic, data sample, followed by the reversal of this process until it converges to the desired complex data distribution. What sets DMs apart from earlier generative models is their execution over iterative time steps, both forward and backward in time and denoted by $t$, as depicted in Figure 1. The forward and backward diffusion processes are distinguished by:
@@ 3 Diffusion Models Basics
3.1 Basics
Consider a dataset of LR-HR image-pairs, $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}^{N}_{i=1}$. For each time step $t$, the random variable $\mathbf{z}_{t}$ describes the current state. The time step $t$ increases incrementally during forward diffusion, whereas it propagates towards 0 during backward diffusion. In literature, there is no clear distinction between $\mathbf{z}_{t}$ in the forward diffusion and $\mathbf{z}_{t}$ in the backward diffusion. During forward diffusion, we assume $\mathbf{z}_{t}\sim q\left(\mathbf{z}_{t}\mid\mathbf{z}_{t-1}\right)$. Conversely, in the backward diffusion, we assume $\mathbf{z}_{t-1}\sim p\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$. We will denote $T$ with $0<t\leq T$ as the maximal time step for finite cases. The initial data distribution ($t=0$) is represented by $\mathbf{z}_{0}\sim q\left(\mathbf{x}\right)$, which is then slowly injected with noise (additive). Vice versa, DMs remove noise therein by running a parameterized model $p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$ in the reverse time direction that approximates the ideal (but unattainable) denoised distribution $p\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$.
The explicit implementation of the forward diffusion $q$ and backward diffusion $p$, approximated by $p_{\theta}$, is influenced by the specific DM in use. Three primary formulations are prevalent in this domain: Denoising Diffusion Probabilistic Models (DDPMs), Score-Based Generative Models (SGMs), and Stochastic Differential Equations (SDEs) [45]. Each of these types will be discussed next.
3.2 Denoising Diffusion Probabilistic Models (DDPMs)

### 06 Feb 2024 v2

@@ Abstract
Diffusion Models (DMs) have disrupted the image Super-Resolution (SR) field and further closed the gap between image quality and human perceptual preferences. They are easy to train and can produce very high-quality samples that exceed the realism of those produced by previous generative methods. Despite their promising results, they also come with new challenges that need further research: high computational demands, comparability, lack of explainability, color shifts, and more. Unfortunately, entry into this field is overwhelming because of the abundance of publications. To address this, we provide a unified recount of the theoretical foundations underlying DMs applied to image SR and offer a detailed analysis that underscores the unique characteristics and methodologies within this domain, distinct from broader existing reviews in the field. This survey articulates a cohesive understanding of DM principles and explores current research avenues, including alternative input domains, conditioning techniques, guidance mechanisms, corruption spaces, and zero-shot learning approaches. By offering a detailed examination of the evolution and current trends in image SR through the lens of DMs, this survey sheds light on the existing challenges and charts potential future directions, aiming to inspire further innovation in this rapidly advancing area.
@@
Image Super-Resolution, Diffusion Models, Generative Models, Survey.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
In the ever-evolving field of computer vision, the task of image Super-Resolution (SR) – enhancing Low-Resolution (LR) images into High-Resolution (HR) counterparts – stands as a longstanding yet still perplexing challenge, owing to its inherently ill-posed nature: Multiple HR reconstructions are plausible for any given LR image, differing in aspects such as brightness and color [1]. The applications span a broad spectrum, from enhancing everyday photography to refining satellite [2] and medical images [3], all driven forward by recent rapid advancements in deep learning.
The advent of Diffusion Models (DMs) marks a significant shift in generative image generation tasks, including image SR, challenging the long-standing dominance of Generative Adversarial Networks (GANs) [4, 5, 6, 7]. Despite the notable achievements of previous generative models, each comes with its set of limitations. For example, the computational demands of Autoregressive models often outweigh their utility in generating HR images [8, 9, 10], while Normalizing Flows (NFs) and Variational Autoencoders (VAEs) struggle to match the quality expectations. GANs, although powerful, necessitate careful regularization and optimization strategies to mitigate issues like optimization instability and mode collapse [11].
In this context, DMs have not only contributed significantly to SR but have also heralded a new epoch in generative AI. Their pivotal roles in applications such as Dall-E and Stable Diffusion demonstrate that DMs have surpassed GANs in various aspects, redefining the generative models landscape [12, 13, 14]. Their capability to generate high-fidelity, detailed images from LR inputs has shown immense promise in SR, closely aligning with the qualitative judgments of human evaluators and offering a novel approach to image quality enhancement [15]. In other words, human raters perceive SR images generated by DMs as more realistic than those generated by GANs.
However, as the volume of research on DMs continues to expand, it is becoming more challenging, particularly for those new to the field, to stay updated on the latest developments. DMs diverge fundamentally from their generative predecessors, posing new challenges while addressing the limitations of earlier models. Amidst this rapid expansion, identifying coherent trends and potential research directions can be challenging. This survey aims to demystify the current landscape of DMs in SR, offering a comprehensive overview that bridges foundational concepts with the forefront of innovation.
This survey builds upon the previous work Hitchhiker’s Guide to Super-Resolution [16], which gives a broad overview of the image SR field in general. We explore the foundational principles of DMs, dive into the challenges specific to SR, and examine the various types of DMs (DDPMs [4], SGMs [7], SDEs [17]), their relationship with other generative models, efficient sampling techniques, and applications in fields demanding high precision, such as medical imaging and remote sensing. Further, we investigate alternative input domains, architectural designs, and conditioning strategies and critically analyze current strengths and weaknesses.
Concluding with a discussion on emerging trends and the potential they hold for reshaping SR and DM development, this survey sets the stage for future research. By offering clarity and direction in the rapidly evolving domain of DMs, we aim to inspire and inform the next wave of research, fostering advancements that continue to push the boundaries of what is possible in image SR with DMs.
@@ I Introduction
section II - Super-Resolution Basics: This section provides fundamental definitions and introduces standard datasets, methods, and metrics for assessing image quality commonly utilized in image SR publications.
section III - Diffusion Models Basics: Introduces the principles and various formulations of DMs, including Denoising Diffusion Probabilistic Models (DDPMs), Score-based Generative Models (SGMs), and Stochastic Differential Equations (SDEs). This section also explores how DMs relate to other generative models.
section IV - Improvements for Diffusion Models: Common practices for enhancing DMs, focusing on efficient sampling techniques and improved likelihood estimation.
section V - Diffusion Models for Image SR: Presents concrete realizations of DMs in SR, explores alternative domains (latent space and wavelet domain), discusses architectural designs and multiple tasks with Null-Space Models, and examines alternative corruption spaces.
section VI - Domain-Specific Applications: DM-based SR applications, namely medical imaging, blind face restoration, atmospheric turbulence in face SR, and remote sensing.
section VII - Discussion and Future Work: Common problems of DMs for image SR and noteworthy research avenues for DMs specific to image SR.
section VIII - Conclusion: Summarizes the survey.
II Image Super-Resolution
Image Super-Resolution (SR) represents a collection of methodologies designed to construct High-Resolution (HR) images from one or more Low-Resolution (LR) images. The domain of SR can be broadly categorized into two areas [16]: Single Image Super-Resolution (SISR) and Multi-Image Super-Resolution (MISR). In the SISR paradigm, a single LR image is processed to yield a single HR image. In contrast, MISR methods utilize multiple LR images to synthesize one or many HR outputs. This section primarily focuses on the SISR facet. It explores relevant datasets, established SR models, and techniques employed to assess image quality.
Given a LR image $\mathbf{x}\in\mathbb{R}^{\bar{w}\times\bar{h}\times c}$, the objective is to produce a corresponding HR image $\mathbf{y}\in\mathbb{R}^{w\times h\times c}$, where $\bar{w}<w$ and $\bar{h}<h$. The relationship is represented through a degradation mapping, defined as
@@ II Image Super-Resolution
where $\mathcal{D}$ is the degradation map $\mathcal{D}:\mathbb{R}^{w\times h\times c}\rightarrow\mathbb{R}^{\bar{w}\times\bar{h}\times c}$ and $\Theta$ signifies the degradation parameters, which can include aspects such as blur type, noise, scaling factor, among others.
The degradation is typically uncertain, posing the challenge for SISR to discern the inverse mapping of $\mathcal{D}$. The inherent complexity of this task arises from the ill-posed nature of predicting $\Theta$, leading to an optimization task aimed at minimizing the discrepancy between the predicted SR image $\mathbf{\hat{y}}$ and the original HR image $\mathbf{y}$:
@@
II-A Datasets
Several datasets, offering a variety of images, resolutions, and content types, are available for SISR and instrumental in effectively training and evaluating SR models. Typically, these datasets consist of LR and HR image pairs. However, certain datasets contain only HR images, with the pairs created by downsampling them using techniques such as bicubic downsampling with anti-aliasing - a default setting for imresize in MATLAB [18]. For natural images, commonly employed datasets include the Berkeley Segmentation Dataset (BSDS300) [19], Flickr2K [20], and the Diverse 2K resolution (DIV2K) dataset [21]. Additionally, datasets such as ImageNet [22] and the Visual Object Classes (VOC2012) [23], which are favored in the Computer Vision (CV) community, are also utilized for SISR tasks. In face SR, datasets featuring high-quality, diverse images of human faces are required. To meet this demand, datasets like CelebA-HQ [24] and Flickr-Faces-HQ (FFHQ) [25] are commonly employed.
II-B SR Models
@@
where $\mathbf{\hat{y}}$ is the predicted HR approximation of the LR image $\mathbf{x}$ and $\theta$ the parameters of $\mathcal{M}$. Consequently, this translates into an optimization objective that minimizes Equation 2, loss function $\mathcal{L}$ between the estimation $\mathbf{\hat{y}}$ and the ground-truth HR image $\mathbf{y}$. The following sections will focus on standard methods to design a SR model, especially conventional deep learning methods, before we examine in detail how diffusion models fulfill the role of a SR model.
II-B1 Traditional Methods
Traditional methods for image SR define a range of methodologies, such as statistical [26], edge-based [27], [28] patch-based [29], [30], prediction-based [31], [32] and sparse representation techniques [33]. They fundamentally rely on image statistics and the information inherent in existing pixels to generate HR images. Despite their utility, a noteworthy drawback of these methods is the potential introduction of noise, blur, and visual artifacts [16].
II-B2 Deep Learning Methods
Image SR has significantly evolved with advancements in Deep Learning (DL) and computational power. In general, DL-based approaches provide notable improvements over traditional methods. Typically, they employ a learning Convolutional Neural Network (CNN) for end-to-end mapping from LR to HR. Initial models, such as SRCNN [34], FSRCNN [35], and ESPCNN [36], utilized simple CNNs of diverse depth and feature maps sizes.
Later models adapted concepts from the broader CV domain into SR models, e.g., ResNet led to SRResNet, where residual information was propagated to successive network layers [37]. Likewise, DenseNet was adapted to image SR[38], e.g., SRDenseNet [39] and ESRGAN [40]. They employ dense blocks, where each layer receives additionally the features generated in all preceding layers. Recursive CNNs that recursively use the same module to learn representations were also inspired by other CV methods for DL-based SR methods in DRCN [41], DRRN [42] and CARN [43]. More recently, attention mechanisms have been incorporated to focus on regions of interest in images, predominantly via the channel and spatial attention mechanisms [16].
@@ II-B2 Deep Learning Methods
II-B3 Pixel-based Methods
Pixel-based techniques utilize loss functions designed to minimize the difference directly in pixel space. Commonly used loss functions are the L1 and L2 losses. Models trained with pixel-based loss functions tend to generate images with a tendency toward blurriness and excessive smoothing [16]. This can result in less high-frequency details and even lead to unrealistic visual appearances. One way to overcome this limitation is to use generative models, such as GANs, flow-based methods, or diffusion models.
II-B4 Generative Adversarial Networks (GANs)
@@
II-B5 Flow-based Methods
@@
II-C Image Quality Assessment (IQA)
@@
II-C1 Peak Signal-to-Noise Ratio (PSNR)
@@
II-C2 Structural Similarity Index (SSIM)
@@
II-C3 Mean Opinion Score (MOS)
The MOS is a subjective measure that leverages human perceptual quality for the evaluation of the generated SR images. Human viewers are shown SR images and asked to rate them with quality scores that are then mapped to numerical values and later averaged. Typically, these range from 1 (bad) to 5 (good) but may vary [15]. While this method is a direct evaluation of human perception, it is more time-consuming and cumbersome to conduct compared to objective metrics. Moreover, due to the highly subjective nature of this metric, it is susceptible to bias.
II-C4 Consistency
@@
II-C5 Learned Perceptual Image Patch Similarity (LPIPS)
@@
III Diffusion Models Basics
Diffusion Models (DMs) have profoundly impacted the realm of generative AI, and many approaches that fall under the umbrella term DM have emerged. At their core, these models share a fundamental principle called diffusion: generate data by gradually introducing noise into an initial, often simplistic, data sample. Next, reverse this process until it converges to the desired complex data distribution. What sets DMs apart from earlier generative models is their execution over iterative time steps, both forward and backward in time and denoted by $t$, as depicted in Figure 1. The forward and backward diffusion processes are distinguished by:
@@ III Diffusion Models Basics
Figure 1: Principle of diffusion models. The forward diffusion adds noise iteratively (red), which translates an image from the image space to the corruption space. The backward diffusion, the iterative refinement process, reverts the process (blue) back to the image space. Shown are three different implementations of diffusion models, namely Denoising Diffusion Probabilistic Models (DDPMs), Score-based Generative Models (SGMs), and Stochastic Differential Equations (SDEs) with their respect formulation of the forward and backward diffusion.
III-A The Core Concept
Consider a dataset of LR-HR image-pairs, $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}^{N}_{i=1}$. For each time step $t$, the random variable $\mathbf{z}_{t}$ describes the current state, which constitutes an image between the image and corruption space. The time step $t$ increases incrementally during forward diffusion, whereas it propagates towards 0 during backward diffusion. In literature, there is no clear distinction between $\mathbf{z}_{t}$ in the forward diffusion and $\mathbf{z}_{t}$ in the backward diffusion. During forward diffusion, we assume $\mathbf{z}_{t}\sim q\left(\mathbf{z}_{t}\mid\mathbf{z}_{t-1}\right)$. Conversely, in the backward diffusion, we assume $\mathbf{z}_{t-1}\sim p\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$.
We will denote $T$ with $0<t\leq T$ as the maximal time step for finite cases. The initial data distribution ($t=0$) is represented by $\mathbf{z}_{0}\sim q\left(\mathbf{x}\right)$, which is then slowly injected with noise (additive). Vice versa, DMs remove noise therein by running a parameterized model $p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$ in the reverse time direction that approximates the ideal (but unattainable) denoised distribution $p\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t}\right)$.
The explicit implementation of the forward diffusion $q$ and backward diffusion $p$, approximated by $p_{\theta}$, is influenced by the specific DM in use. DMs can coarsely be divided into three types: Denoising Diffusion Probabilistic Models (DDPMs), Score-Based Generative Models (SGMs), and Stochastic Differential Equations (SDEs) [48]. Each of these types will be discussed next. The distinctive formulations of the forward and backward diffusion are also comprehensively shown in Figure 1.
III-B Denoising Diffusion Probabilistic Models (DDPMs)